The key is S=C+rC
A watch costing $98 is sold for 156.80. Find the markup rate on the watch.
How do you solve this equation

Answer by stanbon(75887)

S=C+rC
A watch costing $98 is sold for 156.80. Find the markup rate on the watch.
---------------
156.80 = 98 + r*98
156.80 = 98(1+r)
1+r = 1.6
r = 0.60 = 60%
========================
Cheers,
Stan H.